Balloon has a volume of 600-mL at temperature of 300 K. If the temperature of the balloon drops to 250 K, find the volume (in mL)


1
Expert's answer
2021-05-12T06:28:40-0400

V1/T1=V2/T2

0.6 / 300 = V2 / 250

V2 = (0.6 / 300) x 250 = 0.5 L = 500 ml